# A two-phase model of galaxy formation: III. The formation of globular clusters

### v1

@@ 3.5 The fate of sub-clouds and the two channels of globular cluster formation
The star formation threshold density $n_{\rm sf,1}$ obtained in Eq. (44) is sensitive to metallicity. In the high-redshift Universe where pristine inflow dominates the gas accretion, the metallicity of sub-clouds can be as low as $Z\approx 0.02Z_{\odot}$ so that the density at the triggering of fragmentation can reach $n_{\rm snf}$. Thus, the low-metallicity environment naturally gives rise to the formation of globular clusters, regardless of the initial density produced by shocks. On the other hand, for galactic environments where gas is already enriched, sub-clouds with low initial densities can fragment before they reach $n_{\rm snf}$. This will lead to the formation of star clusters under the influence of supernova feedback, making them less dense, less massive, and more fragile to the tidal field of the host galaxy. Such a star cluster is expected to become an open cluster or to disperse and join the diffuse stellar component. In metal-enriched galaxies, therefore, a globular cluster can form only if the initial density of the parent sub-cloud is $\gtrsim n_{\rm snf}$.
@@
Our definition of globular clusters resembles the definition of Pop-I and Pop-II stars, in the sense that Pop-I GCs are metal-rich and Pop-II GCs are metal-poor. Observationally, Pop-I GCs appear red in color, while Pop-II GCs appear blue. Other phenomenological classifications of GCs, such as those based on in- and ex-situ origins, can be derived by integrating the star formation history and merger history of a given system (galaxy/halo).
@@ 3.6 Dynamical evolution
Due to the large sample and dynamic range that we aim to cover, the spatially resolved tidal environment is not available for individual star clusters. Following Chen & Gnedin (2023) and Chen & Gnedin (2024), we adopt a semi-analytical approach, which is calibrated by the high-resolution N-body simulations of Gieles & Gnedin (2023), with modifications according to the SGC properties from our model. For a star cluster with mass $M_{\rm cls}(t)$ moving in an orbit with radius $R_{\rm orb}(t)$ and $V_{\rm orb}(t)$ around the host galaxy, the total mass loss rate can be modeled as
@@
where $M_{\rm cls}(z_{0})=2\times 10^{5}{\rm M_{\odot}}\,M_{\rm cls,5.3}(z_{0})$, $f_{\rm gas}=0.04f_{\rm gas,0.04}$, and $R=0.4$ is the mass loss due to stellar evolution assuming a Chabrier IMF. Note that the change of tidal environment in mergers is difficult to model. We ignore such effects in our modeling and apply Eq. (3.6) to all star clusters, regardless of whether or not they have merged into other galaxies. This is a rough approximation, but may be reasonable given that tidal disruption is more significant at higher $z$, so that most of the mass loss is expected to occur in the galaxy where star clusters were born (e.g. Kruijssen, 2014).
@@
The disruption time scale is proportional to the Hubble time scale, and does not depend on halo mass, indicating that the GCS mass-halo mass relation is not distorted by the dynamical evolution, while the normalization decays if the GCs stop forming. The short disruption time scale at $z=9$ indicates that nearly all GCs formed at such high redshift are eventually disrupted. The relatively long disruption time scale at $z=2$ suggests that massive GCs born with $M_{\rm cls,5.3}\gtrsim 1$ at $z=2$ can survive until $z=0$.
@@ 4 Model predictions and comparison with observational data
The simulation used here is the TNG100-1-Dark run, conducted as a part of the IllustrisTNG project (Pillepich et al., 2018; Nelson et al., 2019). The cosmology adopted by the simulation is consistent with that used in this paper (see §1). The run has a periodic box with a side length of $75\,h^{-1}{\rm{Mpc}}$, $1820^{3}$ dark matter particles, each with a mass of $6.0\times 10^{6}\,h^{-1}{\rm M_{\odot}}$, a Plummer equivalent gravitational softening length varying from $1\,h^{-1}{\rm{kpc}}$ at high $z$ to $0.5\,h^{-1}{\rm{kpc}}$ at low $z$. A total of 100 snapshots spanning from redshift $z=20.0$ to 00 have been saved. Halos are identified using the friends-of-friends (FoF) algorithm with a scaled linking length of $0.2$ (Davis et al., 1985). Subhalos are identified using the Subfind algorithm (Springel et al., 2001; Dolag et al., 2009), and subhalo merger trees are constructed using the SubLink algorithm (Springel, 2005; Boylan-Kolchin et al., 2009; Rodriguez-Gomez et al., 2015). The lower limit for FoF halo mass is about $2\times 10^{8}\,h^{-1}{\rm M_{\odot}}$. The main progenitor of a subhalo is defined as the one with the most massive history among all progenitors (De Lucia & Blaizot, 2007; Rodriguez-Gomez et al., 2015). The central subhalo of a FoF halo is defined as the one with the most massive history among all subhalos within the halo. The main branch of an FoF halo corresponds to the main branch of its central subhalo. The specific run we choose ensures a balance between statistical robustness and resolution in halo assembly histories.
@@ 4.2 Mass and age distributions
Figure 8: Mass functions of GCs. Different panels show the mass functions of GCs in central galaxies with different halo masses at $z=0$. Both in-situ and ex-situ GCs are included. In each panel, grey curve shows the birth-time mass function of all GCs formed in the history, while black curve shows the $z=0$ mass function of GCs survived after the mass loss due to dynamical evolution is incorporated. Red and blue curves show the Pop-I and Pop-II sub-components, respectively, of the black curve. See §4.1 for the details of modeling and sample selection. In the center panel, the histograms show the observed GC mass functions for MW (green) and M31 (purple), compiled by Rodriguez et al. (2023). See §4.2 for a detailed discussion of this figure.
Figure 9: Age distributions of GCs. Different panels show the age distributions of GCs in central galaxies with different halo masses at $z=0$. Both in-situ and ex-situ GCs are included. In each panel, grey curve shows the distribution of the birth time of all GCs formed in the history, while black curve shows that of GCs survived until $z=0$ after the mass loss due to dynamical evolution is incorporated. Red and blue curves show the Pop-I and Pop-II sub-components, respectively, of the black curve. See §4.1 for the details of modeling and sample selection. In the center panel, the histograms show the results of the empirical model of Chen & Gnedin (2024) for MW (green) and M31 (purple). Their results for all and survived GCs are shown by dashed and solid histograms, respectively. See §4.2 for a detailed discussion of this figure.
@@
A key difference between our model and theirs is that we include the Pop-II (metal-poor) channel which elevates GC formation in metal-poor galaxies, mainly low-mass galaxies at high $z$. The consequence is the enhanced GC formation at high redshift, which produces an age distribution peaked at $z\approx 5$, roughly the peak redshift for cosmic bulge formation (Paper-I), instead of at $z\approx 2$, the peak redshift of the cosmic star formation history (e.g. Madau & Dickinson, 2014). Thus, a stronger connection between GCs and bulges is predicted by our model in comparison to other models and early observations. The new GC candidates in the galactic bulge discovered recently by Gaia (Palma et al., 2019; Camargo & Minniti, 2019; Minniti et al., 2021) appears to support the prediction of our model. Metal-poor GCs in present-day central galaxies hosted by halos below the MW-halo mass are dominated by the in-situ component, and thus have systematically older age than the metal-rich counterparts, as seen in the left and center panels. On the other hand, for cluster-size halos, the age distribution of metal-poor GCs in their central galaxies is bimodal, as seen from the right panel of Fig. 9. The additional metal-poor, low-$z$ mode in our model is due to the ex-situ GCs brought in by satellite galaxies merged into the central, and it is consistent with the inference by, e.g., Harris (2023) from observations of bright elliptical galaxies.
@@ 4.3 The size-mass relation
Figure 10: Size-mass relation of GCs. Thick black curve shows the model prediction of the birth-time median size at a given mass by stacking all GCs formed in the histories of central galaxies with $10^{11.5}\leqslant M_{{\rm v},z=0}/(\,h^{-1}{\rm M_{\odot}})<10^{12.5}$. Both in-situ and ex-situ GCs are included in the analysis. Here $R_{\rm cls}$ is defined as the half-mass radius of GC. Shaded areas, from inner to outer, encompass 1, 2 and 3-$\sigma$ ranges, respectively. Red and blue curves show relations of Pop-I and Pop-II GCs, respectively. Star clusters compied by Krumholz et al. (2019) from observations for individual galaxies also shown: Each cross marker represents the median and 1-$\sigma$ range of all observed massive ($M_{*}\gtrsim 10^{4}{\rm M_{\odot}}$) star clusters in a given galaxy; each green circle represents a single YMSC in MW, with the error bar indicating the uncertainty for this star cluster. Brown circles with error bars are running median and 1-$\sigma$ uncertainty for a large sample of YMSCs in the LEGUS survey obtained by Brown & Gnedin (2021). The observed size-mass trend of GC is very weak and the scatter of GC size is large even in a single galaxy. The power-law index of our prediction, due to the adoption of wind feedback, is about $1/5$, as indicated by the black dashed line (see Eq. 3.4), consistent with the observation of Brown & Gnedin (2021) for YMSCs. The present-day size-mass relation of GCs after taking into account the dynamical evolution is expected to be the similar. See §4.3 for a detailed discussion of this figure.
In addition to marginal distributions of individual GC properties, a number of joint distributions have also been shown in the literature to characterize the interconnection among different physical processes. One of such relations is the GC size-mass relation, or more formally the size distribution conditioned on the mass, $p(R_{\rm cls}|M_{\rm cls})$. The importance of this relation can be understood by applying the virial theorem to individual bound star clusters. At a given GC mass, the distribution of the binding energy only depends on the size distribution which is related to the history of dissipation (cooling) and energy injection, as well as to other GC properties through these processes.
@@
Symbols in Fig. 10 show the sizes and masses of massive star clusters compiled by Krumholz et al. (2019) and obtained by Brown & Gnedin (2021) from the HST/LEGUS survey. Note that classifications of star clusters are not clean-cut and may not carry particular physical significance (see discussion in, e.g. §1 of Krumholz et al., 2019). YMSCs traditionally refer to massive bound star clusters with young age ($M_{\rm cls}\gtrsim 10^{4}{\rm M_{\odot}}$, age$<100\,{\rm Myr}$); GCs in the MW are considered old (age $\gtrsim 6\,{\rm Gyr}$; see Fig. 9) but those in M31 are not; star clusters in some other galaxies do not even have a clear classification. Despite of the wide range of age, nearly all the observed sizes and masses of massive clusters fall well within the 2-$\sigma$ range of our model prediction. The amplitude and slope (in logarithmic scale) of the size-mass relation obtained by Brown & Gnedin (2021) for LEGUS YMSCs are very close to our model prediction, thus providing support to our model. Given the large scatter in the observed cluster size at given cluster mass and the significant galaxy-galaxy variation, it is still difficult to use the observed relation to constrain different feedback channels.
@@
Another uncertainty arises from the dynamical evolution of GCs, which can not only change their masses, as modeled in §3.6, but also modifies their sizes. As our model does not track the change of the GC size during the dynamical evolution, it is unclear whether or not the present-day size-mass relation is different from that at the birth time. Some studies have suggested a ‘two-phase’ scenario for the size evolution in the tidal field (e.g. Gieles & Portegies Zwart, 2011): an early ‘expansion-dominated’ phase before the cluster becomes Roche lobe filling, and a later ‘evaporation-dominated’ phase. The starting point and the duration of the two phases thus depend on the initial conditions for cluster formation. Using N-body simulations, Kruijssen et al. (2011, see their figure 2) found that, during the second phase, the evolution paths of individual star clusters in the size-mass plane follows the observed size-mass relation at $z=0$. Thus, if a population of GCs are born on the local size-mass relation, as suggested by our analysis, they are expected to follow the observed relation after the dynamical evolution. Whether this is a coincidence due to simplifications in the simulations or a general result remains to be clarified.

### v2

@@ 3.5 The fate of sub-clouds and the two channels of globular cluster formation
The star formation threshold density $n_{\rm sf,1}$ obtained in Eq. (44) is sensitive to metallicity. In the high-redshift Universe where pristine inflow dominates the gas accretion, the metallicity of sub-clouds can be as low as $Z\approx 0.02\,{\rm Z}_{\odot}$ so that the density at the triggering of fragmentation can reach $n_{\rm snf}$. Thus, the low-metallicity environment naturally gives rise to the formation of globular clusters, regardless of the initial density produced by shocks. On the other hand, for galactic environments where gas is already enriched, sub-clouds with low initial densities can fragment before they reach $n_{\rm snf}$. This will lead to the formation of star clusters under the influence of supernova feedback, making them less dense, less massive, and more fragile to the tidal field of the host galaxy. Such a star cluster is expected to become an open cluster or to disperse and join the diffuse stellar component. In metal-enriched galaxies, therefore, a globular cluster can form only if the initial density of the parent sub-cloud is $\gtrsim n_{\rm snf}$.
@@
Using Eq. (44) for $n_{\rm sf,1}$, the critical density $n_{\rm sf,1}=n_{\rm snf}$ that separates the two types of GCs immediately yields a critical gas-phase metallicty, $Z_{\rm crit}=10^{-1.68}\,{\rm Z}_{\odot}$. Our definition of GCs resembles the definition of Pop-I and Pop-II stars, in the sense that Pop-I GCs are metal-rich and Pop-II GCs are metal-poor. Observationally, Pop-I GCs appear red in color, while Pop-II GCs appear blue. Other phenomenological classifications of GCs, such as those based on in- and ex-situ origins, can be derived by integrating the star formation history and merger history of a given system (galaxy/halo). Our classification of GCs is based on the properties of sub-clouds prior to the star formation. This should not be confused with the Type-I/II GCs classified by the observed stellar populations (e.g. Milone et al., 2017; Simioni et al., 2020) that reflect the star formation and enrichment history of a GC (see §4 of Milone & Marino, 2022, for a review).
@@ 3.6 Dynamical evolution
Due to the large sample and dynamic range that we aim to cover, the spatially resolved tidal environment is not available for individual star clusters. Following Chen & Gnedin (2023) and Chen & Gnedin (2024), we adopt a semi-analytical approach, which was originally introduced by Fall & Zhang (2001), extended by McLaughlin & Fall (2008), and calibrated by the high-resolution N-body simulations of Gieles & Gnedin (2023), with modifications according to the SGC properties from our model. For a star cluster with mass $M_{\rm cls}(t)$ moving in an orbit with radius $R_{\rm orb}(t)$ and $V_{\rm orb}(t)$ around the host galaxy, the total mass loss rate can be modeled as
@@
where $M_{\rm cls}(z_{0})=2\times 10^{5}{\rm M_{\odot}}\,M_{\rm cls,5.3}(z_{0})$, $f_{\rm gas}=0.04f_{\rm gas,0.04}$, and $R$ is the returned fraction of mass (Eq. 17). Note that the change of tidal environment in mergers is difficult to model. We ignore such effects in our modeling and apply Eq. (3.6) to all star clusters, regardless of whether or not they have merged into other galaxies. This is a rough approximation, but may be reasonable given that tidal disruption is more significant at higher $z$, so that most of the mass loss is expected to occur in the galaxy where star clusters were born (e.g. Kruijssen, 2014).
@@
The dependence of $\tau_{\rm tid}$ on $M_{\rm cls}$ has a power-law index of $2/3$ in our fiducial choice, close to the analytically-estimated and observed values, $0.62$ and $0.60$, respectively, reported by Lamers et al. (2005). The dependence of $\tau_{\rm tid}$ on the tidal frequency (Eq. 66) implies a dependence of $\tau_{\rm tid}$ on the ambient gas density, $\tau_{\rm tid}\sim\rho_{\rm gas}^{-1/2}$, consistent with N-body simulations and observations shown by Lamers et al. (2005). The disruption timescale is proportional to the Hubble timescale, and does not explicitly depend on halo mass, indicating that the GCS mass-halo mass relation is not distorted by the dynamical evolution while the normalization decreases as the formation of GCs stops. The short disruption timescale at $z=9$ indicates that nearly all GCs formed at such high redshift are eventually disrupted. The relatively long disruption timescale at $z=2$ suggests that massive GCs born with $M_{\rm cls,5.3}\gtrsim 1$ at $z=2$ can survive until $z=0$. At $z=0$, the disruption timescale is about $3\,{\rm Gyr}$ for a cluster with $u_{\rm r}=0.5$, $f_{\rm gas}=0.04$ and $M_{\rm cls}=10^{4}{\rm M_{\odot}}$, falling into the range found by Lamers et al. (2005) using observational data. Variations of GC disruption among halos due to their differences in GC formation history are allowed in Eqs. (67) and (3.6) through the gas fraction $f_{\rm gas}$ and time variables $t$ and $z_{0}$. Note that we have ignored substructures of galaxies, such as clumps, spiral arms and bars. These should be included in the future with detailed modeling of the galactic structure.
@@ 4 Model predictions and comparison with observational data
The simulation used here is the TNG100-1-Dark run, conducted as a part of the IllustrisTNG project (Pillepich et al., 2018; Nelson et al., 2019). The cosmology adopted by the simulation is consistent with that used in this paper (see §1). The run has a periodic box with a side length of $75\,h^{-1}{\rm{Mpc}}$, $1820^{3}$ dark matter particles, each with a mass of $6.0\times 10^{6}\,h^{-1}{\rm M_{\odot}}$, a Plummer equivalent gravitational softening length varying from $1\,h^{-1}{\rm{kpc}}$ at high $z$ to $0.5\,h^{-1}{\rm{kpc}}$ at low $z$. A total of 100 snapshots spanning from redshift $z=20.0$ to 00 have been saved. Halos are identified using the friends-of-friends (FoF) algorithm with a scaled linking length of $0.2$ (Davis et al., 1985). Subhalos are identified using the Subfind algorithm (Springel et al., 2001; Dolag et al., 2009), and subhalo merger trees are constructed using the SubLink algorithm (Springel, 2005; Boylan-Kolchin et al., 2009; Rodriguez-Gomez et al., 2015). The lower limit for FoF halo mass is about $2\times 10^{8}\,h^{-1}{\rm M_{\odot}}$, and the most massive halo at $z=0$ in the simulation volume has $M_{\rm v}=2.47\times 10^{14}\,h^{-1}{\rm M_{\odot}}$. The main progenitor of a subhalo is defined as the one with the most massive history among all progenitors (De Lucia & Blaizot, 2007; Rodriguez-Gomez et al., 2015). The central subhalo of a FoF halo is defined as the one with the most massive history among all subhalos within the halo. The main branch of an FoF halo corresponds to the main branch of its central subhalo. The specific run we choose ensures a balance between statistical robustness and resolution in halo assembly histories.
@@ 4.2 Mass and age distributions
Figure 9: Mass functions of GCs. Different panels show the mass functions of GCs in central galaxies with different halo masses at $z=0$. Both in-situ and ex-situ GCs are included. In each panel, grey curve shows the birth-time mass function of all GCs formed in the history, while black curve shows the $z=0$ mass function of GCs survived after the mass loss due to dynamical evolution is incorporated. Red and blue curves show the Pop-I and Pop-II sub-components, respectively, of the black curve. See §4.1 for the details of modeling and sample selection. In the center panel, the histograms show the observed GC mass functions for MW (green) and M31 (purple), compiled by Rodriguez et al. (2023). Orange markers show the mass function of star clusters in M51, obtained from the star-cluster catalog of LEGUS (Messa et al., 2018). See §4.2 for a detailed discussion of this figure.
Figure 10: Age distributions of GCs. Different panels show the age distributions of GCs in central galaxies with different halo masses at $z=0$. Both in-situ and ex-situ GCs are included. In each panel, grey curve shows the distribution of the birth time of all GCs formed in the history, while black curve shows that of GCs survived until $z=0$ after the mass loss due to dynamical evolution is incorporated. Red and blue curves show the Pop-I and Pop-II sub-components, respectively, of the black curve. See §4.1 for the details of modeling and sample selection. In the center panel, the histograms show the results of the empirical model of Chen & Gnedin (2024) for MW (green) and M31 (purple). Their results for all and survived GCs are shown by dashed and solid histograms, respectively. Orange markers show the age distribution of star clusters more massive than $10^{5}{\rm M_{\odot}}$ in M51, obtained from the star-cluster catalog of LEGUS (Messa et al., 2018). See §4.2 for a detailed discussion of this figure.
@@
A key difference between our model and theirs is that we include the Pop-II (metal-poor) channel which elevates GC formation in metal-poor galaxies, mainly low-mass galaxies at high $z$. The consequence is the enhanced GC formation at high redshift, which produces an age distribution peaked at $z\approx 5$, roughly the peak redshift for cosmic bulge formation (Paper-I), instead of at $z\approx 2$, the peak redshift of the cosmic star formation history (e.g. Madau & Dickinson, 2014). Thus, a stronger connection between GCs and bulges is predicted by our model in comparison to other models and early observations. The new GC candidates in the galactic bulge discovered recently by Gaia (Palma et al., 2019; Camargo & Minniti, 2019; Minniti et al., 2021) appears to support the prediction of our model. The empirical model of GC formation by Valenzuela et al. (2021, see also ) also supports the presence of a distinct channel in the early Universe. Metal-poor GCs in present-day central galaxies hosted by halos below the MW-halo mass are dominated by the in-situ component, and thus have systematically older age than the metal-rich counterparts, as seen in the left and center panels. On the other hand, for cluster-size halos, the age distribution of metal-poor GCs in their central galaxies is bimodal, as seen from the right panel of Fig. 10. The additional metal-poor, low-$z$ mode in our model is due to the ex-situ GCs brought in by satellite galaxies merged into the central, and it is consistent with the inference by, e.g., Harris (2023) from observations of bright elliptical galaxies.
The difference between MW and M31 in the mass and age distributions of GCs indicates the diversity of GC formation in galaxies with different states and assembly histories. To clearly see this, we consider another galaxy, M51 (NGC 5194), a spiral galaxy interacting with a massive companion, NGC 5195. The stellar mass of M51 is inferred to be $\approx 10^{10.6}{\rm M_{\odot}}$ (Leroy et al., 2008; Wei et al., 2021), slightly smaller than that of MW. We take the star-cluster candidates of the LEGUS survey for M51 (Messa et al., 2018), remove the contamination by selecting only those with machine-learning class 1 or 2, and exclude those with age less than $10\,{\rm Myr}$ to avoid unbound systems. The mass distribution of the star clusters so obtained is shown in the center panel of Fig. 9. We further restrict the sample to massive clusters with $M_{\rm cls}\geqslant 10^{5}{\rm M_{\odot}}$ so that they are age-complete (see Fig. 5 of Messa et al., 2018), and show the age distribution of these clusters in the center panel of Fig. 10. At $M_{\rm cls}\gtrsim 10^{5}{\rm M_{\odot}}$, the cluster mass function of M51 is much lower than those of MW, M31 and the model prediction. The age distribution shows that these massive clusters are mostly born at $z<0.5$ by recent star formation, and none of them are as old as those produced at $z\approx 5$, the peak of the GC-formation activity predicted by our model. Such a situation can be created if M51 forms very late, and is hosted by a halo with very stable dynamical environment in the early stage. The lack of galactic-scale turbulence thus suppresses the formation of old and massive GCs, and large amounts of GCs are produced only later by gas compression through, e.g. disk instabilities and/or external interactions (see the discussion in §3.1).
@@ 4.3 The size-mass relation
Figure 11: Size-mass relation of GCs. Thick black curve shows the model prediction of the birth-time median size at a given mass by stacking all GCs formed in the histories of central galaxies with $10^{11.5}\leqslant M_{{\rm v},z=0}/(\,h^{-1}{\rm M_{\odot}})<10^{12.5}$. Both in-situ and ex-situ GCs are included in the analysis. Here $R_{\rm cls}$ is defined as the half-mass radius of GC. Shaded areas, from inner to outer, encompass 1, 2 and 3-$\sigma$ ranges, respectively. Red and blue curves show relations of Pop-I and Pop-II GCs, respectively. Star clusters compied by Krumholz et al. (2019) from observations for individual galaxies also shown: Each cross marker represents the median and 1-$\sigma$ range of all observed massive ($M_{*}\gtrsim 10^{4}{\rm M_{\odot}}$) star clusters in a given galaxy; each green circle represents a single YMSC in MW, with the error bar indicating the uncertainty for this star cluster. For MW GCs, we compile the catalog of Harris (1996, 2010 edition) and Baumgardt & Hilker (2018), and show the results for all, Pop-I and Pop-II of them by black, red and blue cross markers, respectively. Brown circles with error bars are running median and 1-$\sigma$ uncertainty for a large sample of YMSCs in the LEGUS survey obtained by Brown & Gnedin (2021). The observed size-mass trend of GC is very weak and the scatter of GC size is large even in a single galaxy. The power-law index of our prediction, due to the adoption of wind feedback, is about $1/5$, as indicated by the black dashed line (see Eq. 3.4), consistent with the observation of Brown & Gnedin (2021) for YMSCs. The present-day size-mass relation of GCs is expected to be similar if GCs evolve along their birth-time relation. See §4.3 for a detailed discussion of this figure.
In addition to marginal distributions of individual GC properties, a number of joint distributions have also been shown in the literature to characterize the interconnection among different physical processes. One of such relations is the GC size-mass relation, or more formally the size distribution conditioned on the mass, $p(R_{\rm cls}|M_{\rm cls})$. The importance of this relation can be understood by applying the virial theorem to individual bound star clusters. At a given GC mass, the binding-energy distribution of GCs at the birth-time depends on their stellar density profiles which are related to the histories of gas dissipation (cooling). The subsequent dynamical processes of a GC, such as mass segregation, binary hardening, black-hole heating, core collapse, evaporation and tidal heating, can cause changes in the distribution of stars and exchanges between the cluster and its environment (see Portegies Zwart et al., 2010, for a review). The histories of these processes can thus leave imprints on the size of a GC, and shape the final size-mass relation of GCs.
@@
Symbols in Fig. 11 show the sizes and masses of massive star clusters compiled by Krumholz et al. (2019) and obtained by Brown & Gnedin (2021) from the HST/LEGUS survey. For MW GCs, we take the catalog of Harris (1996, 2010 edition) for metallicity, and incorporate the size and mass measurements from Baumgardt & Hilker (2018). We use a two-component Gaussian mixture model to fit the metallicity distribution of MW GCs, classify them into Pop-I and Pop-II, and show the sizes and masses separately for all, Pop-I and Pop-II GCs using three markers with different colors. Note that classifications of star clusters are not clean-cut and may not carry particular physical significance (see discussion in, e.g. §1 of Krumholz et al., 2019). YMSCs traditionally refer to massive bound star clusters with young age ($M_{\rm cls}\gtrsim 10^{4}{\rm M_{\odot}}$, age$<100\,{\rm Myr}$); GCs in the MW are considered old (age $\gtrsim 6\,{\rm Gyr}$; see Fig. 10) but those in M31 are not; star clusters in some other galaxies do not even have a clear classification. Despite of the wide range of age, nearly all the observed sizes and masses of massive clusters fall well within the 2-$\sigma$ range of our model prediction. The amplitude and slope (in logarithmic scale) of the size-mass relation obtained by Brown & Gnedin (2021) for LEGUS YMSCs are very close to our model prediction, thus providing support to our model. Given the large scatter in the observed cluster size at given cluster mass and the significant galaxy-galaxy variation, it is still difficult to use the observed relation to constrain different feedback channels.
@@
Another uncertainty arises from the dynamical evolution of GCs, which can not only change their masses, as modeled in §3.6, but also modifies their sizes. As our model does not track the change of the GC size during the dynamical evolution, it is unclear whether or not the present-day size-mass relation is different from that at the birth time. Some studies have suggested a ‘two-phase’ scenario for the size evolution in the tidal field (e.g. Gieles & Portegies Zwart, 2011): an early ‘expansion-dominated’ phase before the cluster becomes Roche lobe filling, and a later ‘evaporation-dominated’ phase. The starting point and the duration of the two phases thus depend on the initial conditions for cluster formation. Using N-body simulations, Kruijssen et al. (2011, see their figure 2) found that, during the second phase, the evolution paths of individual star clusters in the size-mass plane follows the observed size-mass relation at $z=0$. Thus, if a population of GCs are born on the local size-mass relation, as suggested by our analysis, they are expected to follow the observed relation after the dynamical evolution. We note that the birth-time size-mass relation of star clusters, as well as other initial conditions for their subsequent dynamical evolution, is set up by a complex interplay of stellar physics, gas magnetohydrodynamics, and radiative transfer, which is not well understood and constrained theoretically and observationally (see §3 of Portegies Zwart et al., 2010, for a review). Our model that star-clusters are born in the observed size-mass relation and evolve along it thus only represents a possible solution. Whether this is a coincidence due to model simplification or a more general result remains to be determined. Taking MW GCs as an example, the sizes and masses of Pop-I and Pop-II GCs have no obvious difference in median, but the sizes of Pop-II GCs appear to be skewed towards larger values, as seen from the error bar. This is potentially due to the older age, and thus the longer time for dynamical evolution, of Pop-II GCs. Alternatively, it may originate from the different tidal environments of the two populations. In such cases, dynamical evolution is required to drive them off the birth-time size-mass relation which lies below the observed one. Arca Sedda et al. (2024, see their figure 2 and §3) used N-body simulations and reported such a possible path for the size-mass evolution: star clusters are initially small and under-filling the Roche lobes, but expand to the observed range of sizes within a timescale of $\lesssim 1\,{\rm Gyr}$ due to internal dynamical effects.